Hi!
How do I solve the linear systems using substitution?
-2x+4y = 2
y= -6x +20
and
y= -4x+8
7x - 6y =-17

Answer by Cromlix(4381)

Hi there,
-2x + 4y = 2......1
y = -6x + 20.......2
Substitute (-6x + 20) for
y in Eq(1)
-2x + 4y = 2......1
-2x + 4(-6 + 20) = 2
-2x - 24x + 80 = 2
Collect like terms;
-2x - 24x = 2 - 80
-26x = -78
x = 3
Substitute x = 3 into
Eq (2)
y = -6x + 20.......2
y = -6(3) + 20
y = -18 + 20
y = 2
{3, 2}
..............
y = -4x + 8.......1
7x - 6y = -17.....2
Substitute (-4x + 8) for
y in Eq (2)
7x - 6y = -17.....2
7x -6(-4x + 8)= -17
7x + 24x - 48 = -17
Collect like terms;
7x + 24x = -17 + 48
31x = 31
x = 1
Substitute x = 1 into
Eq(1)
y = -4x + 8.......1
y = -4(1) + 8
y = -4 + 8
y = 4
{1, 4}
Hope this helps.
:-)